OpenStudy (anonymous):

how would you get the coordinets on a graph from the equation... y=3/4 -3?

OpenStudy (phi):

I assume you mean y= (3/4)*x -3 to get (x,y) pairs, pick a value for x, and then do the arithmetic to find y. Hint: pick easy numbers for x to make the problem easy. For example, with x=1 y = (3/4)*1 -3 = 3/4 - 3 = -9/4 and one pair is (1, -9/4)

OpenStudy (anonymous):

so then i just do that twice for each equation and graph it?

OpenStudy (phi):

yes. easy as pie!

OpenStudy (anonymous):

wait i still dont understand

OpenStudy (anonymous):

oh im sorry its upside down

OpenStudy (phi):

If you use x=4, and x=-4 you will get nicer numbers for y y = (3/4)*4 -3 = 3-3 =0 y = (3/4)* -4 -3 = -3 -3 = -6 so use (4,0) and (-4,-6)